OpenStudy (anonymous):

Add and simplify if possible:

OpenStudy (anonymous):

\[x-7 \over 2 \] + \[x + 5 \over 2\]

OpenStudy (chaise):

I assume you just want the answer. x-7 + x+5=2x+2 (2x+2)/2 is the answer.

OpenStudy (anonymous):

\[x-7+x+5=2x-2\] so put it all over 2 and get \[\frac{2x-2}{2}=x-1\]

OpenStudy (chaise):

You can then factor out a 2 ^